INTRODUCTION TO THE NON-LINEAR TRACE THEORY AND GENERALISED BOUNDARY VALUE PROBLEMS

LAURENT VÉRON

Abstract. Lectures given by the author at MSRI during the fall of 1997, during the Stochastic Analysis program.

Section I: The linear Dirichlet problem Section II: The non-linear Dirichlet problem Section III: The non-linear trace